### v4

@@ 2.2 Unique Matrix Separation
Here, the expectation is conditional on all the regressors $(X_{1},\ldots,X_{K})$, and also implicitly on all the parameters $\beta_{0}$ and $\Gamma_{0}$, because those are treated as non-random.191919 $\bar{\beta}_{\psi}$ can be viewed as a population version of $\widehat{\beta}_{\psi}$ for an appropriately defined population distribution of $Y$ conditional on $X$. But independent of this interpretation, $\bar{\beta}_{\psi}$ is a convenient tool for discussing the necessary non-collinearity condition on the regressors without requiring asymptotic analysis.
@@
[…]
@@
Providing such global identification conditions for models with low-rank regressors and unknown $R_{0}$ is a new contribution to the interactive fixed effects literature.212121 If the model would not have any idiosyncratic errors (i.e. $E=0$), then $Y-\beta\cdot X=(\beta_{0}-\beta)\cdot X+\Gamma_{0}$, and a natural solution to this identification problem would be to choose $\beta$ as the solution to the rank minimization problem $\min_{\beta\in\mathbb{R}^{K}}\;{\rm rank}\left(Y-\beta\cdot X\right),$ where at the true parameters we have ${\rm rank}\left(Y-\beta_{0}\cdot X\right)={\rm rank}(\Gamma_{0})=R_{0}$, that is, we are minimizing the number of factors required to describe the data. However, once idiosyncratic errors $E$ are present, then this rank minimization does not work, because $Y-\beta\cdot X$ is of large rank for all $\beta$. Our approach here is similar to the “Identification via a Strict Convex Penalty” proposed in Chen and Pouzo (2012).
@@ 3.1 Consistency Results for Low-Rank Regressors
The theorem shows that both $\widehat{\beta}_{*}$ and $\widehat{\beta}_{\psi}$, for $\psi=\psi_{NT}=O\left(1/{\sqrt{\min(N,T)}}\right)$, converge to $\beta_{0}$ at a rate of at least $\sqrt{\min(N,T)}$. The proof of the theorem is provided in the appendix, and is a relatively easy generalization of the proof of Proposition 1. This is because the assumption that all the regressors $X_{k}$ are low-rank allows to easily decouple the contribution of the high-rank matrix $E$ and the low-rank matrix $\beta\cdot X+\Gamma$ to the penalized objective function $Q_{\psi}(\beta)$. However, dealing with the contribution of the idiosyncratic errors $E$ becomes more complicated once high-rank regressors are present, as will be explained in the following.
@@ 3.2.1 Consistency of $\widehat{\beta}_{\psi}$ and $\widehat{\Gamma}_{\psi}$
Let =$\mathbb{C}=\left\{\Theta\in\mathbb{R}^{N\times T}\;|\;\|\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\|_{1}\leq 3\|\Theta-\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\|_{1}\right\}.$ We assume that there exists $\mu>0$, independent from $N$ and $T$, such that for any $\theta\in\mathbb{R}^{NT}$ with ${\rm mat}(\theta)\in\mathbb{C}$ we have $\theta^{\prime}\mathbf{M}_{x}\theta\geq\mu\,\theta^{\prime}\theta$, for all $N$, $T$.
@@
Notice that for $R_{0}=0$ we have $\mathbf{M}_{\lambda_{0}}=\mathbb{I}_{N}$ and $\mathbf{M}_{f_{0}}=\mathbb{I}_{N}$, and therefor =$\mathbb{C}=\{0_{N\times T}\}$, implying that Assumption 1 is trivially satisfied for any $\mu>0$.
@@ 3.2.2 Consistency of $\widehat{\beta}_{*}$
Here, we establish consistency of the nuclear norm minimization estimator $\widehat{\beta}_{*}$ for high-rank regressors. For simplicity we only discuss the case of a single regressor ($K=1$) in the main text, and we simply write $X$ for the $N\times T$ regressor matrix $X_{1}$ in this subsection. The general case of multiple regressors ($K>1$) is discussed in Appendix B.6.
Remember that $\widehat{\beta}_{*}$ is the minimizer of the objective function $\|Y-\beta\cdot X\|_{1}=\|E+(\beta_{0}-\beta)X+\Gamma_{0}\|_{1}=\sum_{r}s_{r}\left(E+(\beta_{0}-\beta)X+\Gamma_{0}\right)$. Asymptotically separating the contribution of the low-rank matrix $\Gamma_{0}$ to the singular values of the sum $E+(\beta_{0}-\beta)X+\Gamma_{0}$ is possible under a strong factor assumption.222222 In Moon and Weidner (2015, 2017) we use the perturbation theory of linear operator to do exactly that. However, characterizing the singular values of the sum of two high-rank matrices $E+(\beta_{0}-\beta)X$ requires results from random matrix theory that are usually only shown under relatively strong assumptions on the distribution of the matrix entries. We therefore first provide a theorem under high-level assumptions, and afterwards discuss how to verify those assumptions using results from random matrix theory. We write SVD for “singular value decomposition” in the following.
@@ 4 Post Nuclear Norm Regularized Estimation
    For $s=0$ set $\widehat{\beta}^{(s)}=\widehat{\beta}_{\psi}$ (or $=\widehat{\beta}_{*}$), the preliminary consistent estimate for $\beta_{0}$.
@@ 5 Implementation and Monte Carlo Simulations
Table 1: Monte Carlo results based on 1000 repetitions for the design specified in display (24). Reported are the bias and standard deviation for the pooled OLS estimator (POLS), the least squares estimator with $R_{0}=2$ factors (LS), the nuclear norm minimizing estimator $\widehat{\beta}_{*}$ (NNmin), the nuclear norm penalized estimator with $\psi=\widehat{\psi}$ (NNpen), the post estimator $\widehat{\beta}^{(s)}$ for $s=1,2,3$ iterations and using $R=\widehat{R}$ factors (POST$(s)$), and the alternative bias correction method (see the appendix) using $R=\widehat{R}$ factors and $s=1,2,3,4,5$ iterations.
@@
6 Extension to Single Index Models
Assumption 2.
  • (i)
  • (ii)
  • (iii)
  • (iv)
    $\partial_{z}\overline{m}_{it}(z_{it}^{0})=0$, for all $i,t$.
  • (a)
  • (b)
  • (c)
Theorem 5.
  • (i)
    $\|\Gamma_{0}\|_{1}=O(\sqrt{NT})$.
  • (ii)
  • (iii)
Then we have $\widehat{\beta}_{\psi}-\beta_{0}=O_{P}(\psi^{1/2})$.
The proof of Theorem 5 is presented in the appendix, where we also discuss how the result could in principle be extended to $K>1$ regressors, which requires some additional technical restrictions. Notice also that the convergence rate of $\psi^{1/2}$ in Theorem 5 is different from the convergence rate $\psi$ obtained in Section 3, but this is likely an artifact of our proof strategy for Theorem 5. Finally, the analog of the nuclear-norm minimizing estimator $\widehat{\beta}_{*}$ to non-linear models is given by $\lim_{\psi\rightarrow 0}\widehat{\beta}_{\psi}$ (limit for fixed $N,T$), but we do not provide results for that limiting estimator here. The goal of this section was not to fully discuss the non-linear case, but to highlight the potential of the nuclear norm penalization approach beyond the linear model that is main focus of this paper.
@@ 7 Conclusions
In this paper we analyze two new estimation methods for interactive fixed effect panel regressions that are based on convex objective functions: (i) nuclear norm penalized estimation, and (ii) nuclear norm minimizing estimation. The resulting estimators can also be applied in situations where the LS estimator may not be consistent, in particular when low-rank regressors are present and the true number of factors is unknown. We provide consistency and convergence rate results for the new estimators of the regression coefficients, and we show how to use them as a preliminary estimator to achieve asymptotic equivalence to the local version of the LS estimator. We have focused on the linear model with homogenous coefficients, which is a natural starting point to understand the usefulness of nuclear norm penalization approach for panel regression models, but there are several ongoing extensions, including developing a unified method to deal with non-linear models, heterogeneous coefficients, treatment effect estimation, nonparametric sieve estimation, and high-dimensional regressors, see Section 6 above, and also Athey, Bayati, Doudchenko, Imbens, and Khosravi (2021) and Chernozhukov, Hansen, Liao, and Zhu (2018).
@@ A.1 An Example of a Non-convex LS Profile Objective Function
[…]
@@ A.2 Alternative Bias Correction
Appendix B Supplementary Appendix
B.1 Proofs for Section 2.1
@@ Appendix B Supplementary Appendix
Before presenting the next lemma it is useful to introduce some further notation. For $\beta\in\mathbb{R}^{K}$ let $\Delta\beta:=\beta-\beta_{0}$. Let $\lambda_{X}$ be an $N\times R_{\rm c}$ matrix such that the column span of $\lambda_{X}$ equals the column span of the $N\times TK$ matrix $[X_{1},\ldots,X_{K}]$. Analogously, let $f_{X}$ be an $T\times R_{\rm r}$ matrix such that the column span of $f_{X}$ equals the column span of the $T\times NK$ matrix $[X_{1}^{\prime},\ldots,X_{K}^{\prime}]$.
@@
From this and (S.8) with $\alpha=\Delta\bar{\beta}_{\psi}$ we obtain for any $\psi>0$ that262626 The bound (S.10) is sufficient for our purposes since we ultimately consider the limit $\psi\rightarrow 0$ here, but for a fixed value of $\psi$ (and $N,T$) this bound is potentially very crude if high-rank regressors $X_{k}$ are present. From Lemma S.4 one could then obtain a sharper bound on $\bar{\beta}_{\psi}-\beta_{0}$ by not using part (i) of Lemma S.2 to simplify $g_{\psi}\left[\left(\mathbf{M}_{\lambda_{0}}(\Delta\bar{\beta}_{\psi}\cdot X)\mathbf{M}_{f_{0}}\right)/\sqrt{NT}\right]$.
@@
00 0 $\geq Q_{\psi}(\Gamma_{0}+\widehat{\Theta}_{\psi})-Q_{\psi}(\Gamma_{0})$
@@
00 0 $\geq L(\Gamma_{0}+\widehat{\Theta}_{\psi})-L(\Gamma_{0})+\frac{\psi}{\sqrt{NT}}\left(\|\Gamma_{0}+\widehat{\Theta}_{\psi}\|_{1}-\|\Gamma_{0}\|_{1}\right)$
@@
$\widehat{\beta}_{\psi}-\beta_{0}$ ::=:= $\widehat{\beta}({\widehat{\Gamma}}_{\psi})-\beta_{0}=\left(\frac{1}{NT}x^{\prime}x\right)^{-1}\left(\frac{1}{NT}x^{\prime}e-\frac{1}{NT}x^{\prime}(\widehat{\gamma}_{\psi}-\gamma_{0})\right).$
@@
00 0 $\geq Q(\widehat{\Gamma}_{\psi})-Q(\Gamma_{0})$
@@
B.5 Sufficient Conditions for Restricted Strong Convexity

### v5

@@ 2.2 Unique Matrix Separation
[…]
@@
The proof is given in the appendix. The proposition considers fixed $N$, $T$, with only $\psi\rightarrow 0$. [^20] Display (S.10) in the appendix provides a bound on $\|\bar{\beta}_{\psi}-\beta_{0}\|$ for finite $\psi$, but the limit $\psi\rightarrow 0$ is what matters most to us, because that limit allows to identify $\beta_{0}$. The statement of the proposition implies that $\lim_{\psi\rightarrow 0}\bar{\beta}_{\psi}=\beta_{0}$. Thus, the proposition provides conditions under which the nuclear norm regularization approach identifies the true parameter $\beta_{0}$. The proposition does not restrict the rank of the regressors, so the result is applicable to both low-rank and high-rank regressors. The assumption $\mathbb{E}(\left.E_{it}\,\right|\,X)=0$ requires strict exogeneity of all regressors, but we will allow for pre-determined regressors in consistency results of Section 3.2 below.
@@
Providing such global identification conditions for models with low-rank regressors and unknown $R_{0}$ is a new contribution to the interactive fixed effects literature. [^21] If the model would not have any idiosyncratic errors (i.e. $E=0$), then $Y-\beta\cdot X=(\beta_{0}-\beta)\cdot X+\Gamma_{0}$, and a natural solution to this identification problem would be to choose $\beta$ as the solution to the rank minimization problem $\min_{\beta\in\mathbb{R}^{K}}\;{\rm rank}\left(Y-\beta\cdot X\right),$ where at the true parameters we have ${\rm rank}\left(Y-\beta_{0}\cdot X\right)={\rm rank}(\Gamma_{0})=R_{0}$, that is, we are minimizing the number of factors required to describe the data. However, once idiosyncratic errors $E$ are present, then this rank minimization does not work, because $Y-\beta\cdot X$ is of large rank for all $\beta$. Our approach here is similar to the “Identification via a Strict Convex Penalty” proposed in Chen and Pouzo (2012). As noted above, the proof of Proposition 1 applies to the population objective function $\bar{\beta}_{\psi}$ for fixed $N$ and $T$. The consistency results for the sample estimators $\widehat{\beta}_{\psi}$ and $\widehat{\beta}_{*}$ in Section 3 are established using different asymptotic arguments, in particular relying on Lemma S.3 in the supplementary appendix rather than on Proposition 1 directly.
@@ 3.1 Consistency Results for Low-Rank Regressors
The theorem shows that both $\widehat{\beta}_{*}$ and $\widehat{\beta}_{\psi}$, for $\psi=\psi_{NT}=O\left(1/{\sqrt{\min(N,T)}}\right)$, converge to $\beta_{0}$ at a rate of at least $\sqrt{\min(N,T)}$. The proof of the theorem is provided in the appendix, and is a relatively easy generalization of the proof of Proposition 1.
We note that condition (18) requires a lower bound of order $c>0$ (a positive constant), whereas condition (17) in Proposition 1 only requires a strict inequality. This difference reflects the distinction between a population identification condition (Proposition 1, for fixed $N$ and $T$) and a sample condition for asymptotic analysis (Theorem 1, as $N,T\rightarrow\infty$). The nuclear norms appearing in (18) are of order $\sqrt{NT}$, so requiring their difference to be bounded below by a positive constant $c$ uniformly over all sample sizes ensures the condition does not become slack asymptotically.
Returning to Example 1, the DGP with low-rank regressor $X_{1}=vw^{\prime}$ satisfies condition (18) when the vectors $v$ and $w$ are sufficiently different from the factor loadings $\lambda_{0}$ and factors $f_{0}$. Specifically, if $v$ is not well-approximated by linear combinations of the columns of $\lambda_{0}$, and similarly $w$ is not well-approximated by linear combinations of the columns of $f_{0}$, then the projection $\mathbf{M}_{\lambda_{0}}X_{1}\mathbf{M}_{f_{0}}=(\mathbf{M}_{\lambda_{0}}v)(\mathbf{M}_{f_{0}}w)^{\prime}$ retains substantial norm, while $\mathbf{P}_{\lambda_{0}}X_{1}\mathbf{P}_{f_{0}}=(\mathbf{P}_{\lambda_{0}}v)(\mathbf{P}_{f_{0}}w)^{\prime}$ remains small. In the extreme case where $v$ lies in the column space of $\lambda_{0}$ or $w$ lies in the column space of $f_{0}$, condition (18) fails and $\beta_{0}$ is not identified, consistent with the non-identification illustrated in Example 1.
@@ 3.2.1 Consistency of $\widehat{\beta}_{\psi}$ and $\widehat{\Gamma}_{\psi}$
Let $\mathbb{C}=\left\{\Theta\in\mathbb{R}^{N\times T}\;|\;\|\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\|_{1}\leq 3\|\Theta-\mathbf{M}_{\lambda_{0}}\Theta\mathbf{M}_{f_{0}}\|_{1}\right\}.$ We assume that there exists $\mu>0$, independent from $N$ and $T$, such that for any $\theta\in\mathbb{R}^{NT}$ with ${\rm mat}(\theta)\in\mathbb{C}$ we have $\theta^{\prime}\mathbf{M}_{x}\theta\geq\mu\,\theta^{\prime}\theta$, for all $N$, $T$.
@@
Notice that for $R_{0}=0$ we have $\mathbf{M}_{\lambda_{0}}=\mathbb{I}_{N}$ and $\mathbf{M}_{f_{0}}=\mathbb{I}_{N}$, and therefor $\mathbb{C}=\{0_{N\times T}\}$, implying that Assumption 1 is trivially satisfied for any $\mu>0$.
@@ 3.2.2 Consistency of $\widehat{\beta}_{*}$
Here, we establish consistency of the nuclear norm minimization estimator $\widehat{\beta}_{*}$ for high-rank regressors. For simplicity we only discuss the case of a single regressor ($K=1$) in the main text, and we simply write $X$ for the $N\times T$ regressor matrix $X_{1}$ in this subsection. The general case of multiple regressors ($K>1$) is discussed in Appendix C.6.
Remember that $\widehat{\beta}_{*}$ is the minimizer of the objective function $\|Y-\beta\cdot X\|_{1}=\|E+(\beta_{0}-\beta)X+\Gamma_{0}\|_{1}=\sum_{r}s_{r}\left(E+(\beta_{0}-\beta)X+\Gamma_{0}\right)$. Asymptotically separating the contribution of the low-rank matrix $\Gamma_{0}$ to the singular values of the sum $E+(\beta_{0}-\beta)X+\Gamma_{0}$ is possible under a strong factor assumption. [^22] In Moon and Weidner (2015, 2017) we use the perturbation theory of linear operator to do exactly that. However, characterizing the singular values of the sum of two high-rank matrices $E+(\beta_{0}-\beta)X$ requires results from random matrix theory that are usually only shown under relatively strong assumptions on the distribution of the matrix entries. We therefore first provide a theorem under high-level assumptions, and afterwards discuss how to verify those assumptions using results from random matrix theory. We write SVD for “singular value decomposition” in the following.
@@
Remark 1.
Several identification-type conditions appear in this paper, and we briefly summarize their relationships here:
  • Condition (17) in Proposition 1 is a population identification condition for fixed $N$ and $T$. It applies to both low-rank and high-rank regressors and requires that the part of any linear combination of regressors not explained by the factor structure dominates (in nuclear norm) the part that can be explained.
  • Condition (18) in Theorem 1 is the sample analog of (17) for low-rank regressors. The $c\sqrt{NT}$ scaling ensures the condition holds uniformly as $N,T\rightarrow\infty$, reflecting the order of magnitude of nuclear norms in the asymptotic analysis.
  • Assumption 1 (Restricted Strong Convexity) applies when regressors are high-rank. It requires that the quadratic form $\theta^{\prime}\mathbf{M}_{x}\theta$ is bounded below on the cone $\mathbb{C}$, which contains matrices close to the factor structure.
  • The conditions in Theorem 3 provide sufficient primitive conditions for consistency of $\widehat{\beta}_{*}$ when regressors are high-rank. In particular, conditions (iv) and (v) ensure that $X$ is sufficiently different from both the factor structure and the idiosyncratic errors.
All these conditions share a common intuition: the regressors must be sufficiently “different” from the factor structure $\Gamma_{0}$ (for low-rank $X$) or from the idiosyncratic errors $E$ (for high-rank $X$) to ensure that $\beta_{0}$ is identified.
We note that our conditions are stated in stochastic form (holding with probability approaching one or for all realizations). In related work, such as Belloni, Chen, Madrid Padilla, and Wang (2023) and Feng (2024), analogous conditions are sometimes stated in expectation form. Relaxing our conditions to expectation form would require verifying that appropriate concentration inequalities hold for the nuclear norm quantities involved, which we have not pursued here. Our stochastic form is standard in this setting and sufficient for our proofs.
@@ 4 Post Nuclear Norm Regularized Estimation
    For $s=0$ set $\widehat{\beta}^{(s)}=\widehat{\beta}_{\psi}$ (or $\widehat{\beta}_{*}$), the preliminary consistent estimate for $\beta_{0}$.
@@
Remark 2.
Step 3 minimizes over $(\beta,g,h)$ jointly rather than using the simpler regression of $Y-\widehat{\lambda}^{(s+1)}\widehat{f}^{(s+1)\prime}$ on $X$. The key advantage of this joint minimization is that it closely approximates a Newton-Raphson step for minimizing the profile objective function $L_{R_{0}}(\beta)$. As shown in Theorem 4, this yields convergence similar to Newton-Raphson: the distance to the local LS estimator shrinks rapidly with each iteration, allowing the procedure to be stopped after just a few steps while still achieving asymptotic equivalence to $\widehat{\beta}^{\rm\,local}_{{\rm LS},R_{0}}$.
@@ 5 Implementation and Monte Carlo Simulations
Table 1: Monte Carlo results based on 1000 repetitions for the design specified in display (24). Reported are the bias and standard deviation for the pooled OLS estimator (POLS), the least squares estimator with $R_{0}=2$ factors (LS), the nuclear norm minimizing estimator $\widehat{\beta}_{*}$ (NNmin), the nuclear norm penalized estimator with $\psi=\widehat{\psi}$ (NNpen), the post estimator $\widehat{\beta}^{(s)}$ for $s=1,2,3$ iterations and using $R=\widehat{R}$ factors (POST(s)(s)), and the alternative bias correction method (see the appendix) using $R=\widehat{R}$ factors and $s=1,2,3,4,5$ iterations.
@@ 6 Conclusions
In this paper we analyze two new estimation methods for interactive fixed effect panel regressions that are based on convex objective functions: (i) nuclear norm penalized estimation, and (ii) nuclear norm minimizing estimation. The resulting estimators can also be applied in situations where the LS estimator may not be consistent, in particular when low-rank regressors are present and the true number of factors is unknown. We provide consistency and convergence rate results for the new estimators of the regression coefficients, and we show how to use them as a preliminary estimator to achieve asymptotic equivalence to the local version of the LS estimator. We have focused on the linear model with homogenous coefficients, which is a natural starting point to understand the usefulness of nuclear norm penalization approach for panel regression models, but there are several ongoing extensions, including developing a unified method to deal with non-linear models, heterogeneous coefficients, treatment effect estimation, nonparametric sieve estimation, and high-dimensional regressors — see Athey, Bayati, Doudchenko, Imbens, and Khosravi (2021) and Chernozhukov, Hansen, Liao, and Zhu (2023). In Appendix B we briefly discuss how the nuclear norm regularization approach extends to nonlinear panel models, including maximum likelihood, weighted least squares, and quantile regression settings. We note that sharper convergence rates than those presented in the appendix have been established for specific nonlinear models in the subsequent literature, including Belloni, Chen, Madrid Padilla, and Wang (2023), Wang, Su, and Zhang (2022), and Feng (2024) for quantile regression.
@@ A.1 An Example of a Non-convex LS Profile Objective Function
and $\lambda_{x,i}\sim iid\,\,2\chi^{2}(1)$, $f_{x,t}\sim iid\,\,2\chi^{2}(1)$, $E_{x,it},E_{it}\sim i.i.d.\,{\cal N}(0,1)$, and $\{\lambda_{0,i}\}$, $\{f_{0,t}\}$, $\{\lambda_{x,i}\}$, $\{f_{x,t}\}$, $\{E_{x,it}\},\{E_{it}\}$ are all independent of each other. For $(N,T)=(200,200)$, we generate the panel data for $(Y_{it},X_{it})$, and plot the LS objective function (3) in Figure 1, which is discussed in the main text.
@@ A.2 Alternative Bias Correction
Appendix B Extension to Nonlinear Panel Data Models
This appendix extends the nuclear norm regularization approach to nonlinear panel data models. The results here are less sharp than those in the main text for linear models; see Remark 3 below and the references cited therein for sharper results in specific nonlinear settings.
Assumption 2.
  • (i)
  • (ii)
  • (iii)
  • (iv)
    $\partial_{z}\overline{m}_{it}(z_{it}^{0})=0$, for all $i,t$.
  • (a)
  • (b)
  • (c)
Theorem A.1.
  • (i)
    $\|\Gamma_{0}\|_{1}=O(\sqrt{NT})$.
  • (ii)
  • (iii)
Then we have $\widehat{\beta}_{\psi}-\beta_{0}=O_{P}(\psi^{1/2})$.
Remark 3.
The convergence rate $O_{P}(\psi^{1/2})$ in Theorem A.1 is slower than the rate $O_{P}(\psi)$ obtained for the linear model in Section 3. This slower rate is likely an artifact of our proof strategy. For specific nonlinear models, sharper results have been established in the subsequent literature. For example, Belloni, Chen, Madrid Padilla, and Wang (2023) and Feng (2024) show that nuclear norm penalized estimators for quantile regression can achieve the same $O_{P}(\psi)$ rate as in the linear case.
The proof of Theorem A.1 is presented in the appendix, where we also discuss how the result could in principle be extended to $K>1$ regressors, which requires some additional technical restrictions. Notice also that the convergence rate of $\psi^{1/2}$ in Theorem A.1 is different from the convergence rate $\psi$ obtained in Section 3, but this is likely an artifact of our proof strategy for Theorem A.1. Finally, the analog of the nuclear-norm minimizing estimator $\widehat{\beta}_{*}$ to non-linear models is given by $\lim_{\psi\rightarrow 0}\widehat{\beta}_{\psi}$ (limit for fixed $N,T$), but we do not provide results for that limiting estimator here. The goal of this section was not to fully discuss the non-linear case, but to highlight the potential of the nuclear norm penalization approach beyond the linear model that is main focus of this paper.
Appendix C Supplementary Appendix
C.1 Proofs for Section 2.1
@@ Appendix C Supplementary Appendix
Before presenting the next lemma it is useful to introduce some further notation. For $\beta\in\mathbb{R}^{K}$ let $\Delta\beta:=\beta-\beta_{0}$. Let $\lambda_{X}$ be an $N\times R_{\rm c}$ matrix such that the column span of $\lambda_{X}$ equals the column span of the $N\times TK$ matrix $X_{1},\ldots,X_{K}]$. Analogously, let $f_{X}$ be an $T\times R_{\rm r}$ matrix such that the column span of $f_{X}$ equals the column span of the $T\times NK$ matrix $X_{1}^{\prime},\ldots,X_{K}^{\prime}]$.
@@
From this and (S.8) with $\alpha=\Delta\bar{\beta}_{\psi}$ we obtain for any $\psi>0$ that [^26] The bound (S.10) is sufficient for our purposes since we ultimately consider the limit $\psi\rightarrow 0$ here, but for a fixed value of $\psi$ (and $N,T$) this bound is potentially very crude if high-rank regressors $X_{k}$ are present. From Lemma S.4 one could then obtain a sharper bound on $\bar{\beta}_{\psi}-\beta_{0}$ by not using part (i) of Lemma S.2 to simplify $g_{\psi}\left[\left(\mathbf{M}_{\lambda_{0}}(\Delta\bar{\beta}_{\psi}\cdot X)\mathbf{M}_{f_{0}}\right)/\sqrt{NT}\right]$.
@@
$0$ $\geq Q_{\psi}(\Gamma_{0}+\widehat{\Theta}_{\psi})-Q_{\psi}(\Gamma_{0})$
@@
$0$ $\geq L(\Gamma_{0}+\widehat{\Theta}_{\psi})-L(\Gamma_{0})+\frac{\psi}{\sqrt{NT}}\left(\|\Gamma_{0}+\widehat{\Theta}_{\psi}\|_{1}-\|\Gamma_{0}\|_{1}\right)$
@@
$\widehat{\beta}_{\psi}-\beta_{0}$ $:=$ $\widehat{\beta}({\widehat{\Gamma}}_{\psi})-\beta_{0}=\left(\frac{1}{NT}x^{\prime}x\right)^{-1}\left(\frac{1}{NT}x^{\prime}e-\frac{1}{NT}x^{\prime}(\widehat{\gamma}_{\psi}-\gamma_{0})\right).$
@@
$0$ $\geq Q(\widehat{\Gamma}_{\psi})-Q(\Gamma_{0})$
@@
C.5 Sufficient Conditions for Restricted Strong Convexity